OpenStudy (anonymous):

What are the x and y intercepts of the quadratic equation y= 2x^2 -8x + 6? 1) x: (-3,0) and (-1,0) y: (0,3) 2) x: (-3,0) and (-1,0) y: (0,6) 3) x: (-3,0) and (-1,0) y: (0,3) 4) x: (3,0) and (1,0) y: (0,6)

OpenStudy (anonymous):

4

OpenStudy (anonymous):

y - intercept: x = 0, so y = 6

OpenStudy (anonymous):

x - intercept, set \[2x^2-8x+6=0\] solve for x

OpenStudy (anonymous):

first divide to 2 to get \[x^2-4x+3=0\] then \[(x-3)(x-1)=0\] so \[x-3=0\] \[x=3\] or \[x-1=0\] \[x=1\]

OpenStudy (anonymous):

number 4 is right

OpenStudy (anonymous):

to find x-intercept, find the solution to 2x^2 -8x + 6=0 So the x-intercepts are 3 and 1 and to find the y-intercept, set all the x values to zero and you have 6.